OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 5x - y = -6 10x + 4y = -6

OpenStudy (mathwizzard3):

Do you know what the substitution method is?

OpenStudy (anonymous):

i just need a quick answer...

OpenStudy (mathwizzard3):

Oh okay :) Let me see

OpenStudy (mathwizzard3):

Do you just need the variables?

OpenStudy (anonymous):

i need the new equation.

OpenStudy (anonymous):

A. 10x + 4(-5x - 6) = -6 B. 10x + 4(5x + 6) = -6 C. 10(-5x - 6) + 4y = -6 D. 10(5x + 6) + 4y = -6

OpenStudy (mathwizzard3):

oh!

OpenStudy (mathwizzard3):

B

OpenStudy (anonymous):

ty : )

OpenStudy (mathwizzard3):

Hope that helps! :D np